# {centering}Tests of restricted Quantum Focusing and a universal CFT bound

### v1

@@ 3 The rQFC and QFC in JT gravity coupled to a QFT
$\left(\frac{\Sigma}{L}\right)^{\delta_{0}-d-1}\gtrsim\left(\frac{\ell_{S}}{L}\right)^{d-2}.$ (4.37)
It is easy to see that this combination of constraints in turn implies:
$\delta_{0}\equiv\text{min}_{j}\delta_{j,n=1}\leq 2(d-1).$ (4.38)
The constraint (4.38) is a necessary condition implied by the rQFC, and, to our knowledge, it is a novel constraint on the CFT spectrum. Note that the bound (4.38) can be equivalently formulated in terms of the singularity structure of two coinciding $\mathcal{E}$ operators, namely that:
$\lim_{|\tilde{y}|\to 0}\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\neq 0,$ (4.39)
or that the $\mathcal{E}\times\mathcal{E}$ OPE is either singular or a non-zero constant.
Let us now investigate this bound in a few simple cases. First, in the case of free theories, the leading Regge trajectory ($n=1$) satisfies:
$\Delta^{\text{even}}_{j=0,1}(J)=d-2+J.$ (4.40)
Therefore, from Eq. (4.32) we have $\delta_{j=0,n=1}=d$, which satisfies the bound (4.38) with room to spare in $d\geq 3$. Since we expect that small interactions would only change the Regge trajectories by a small amount, this suggests that the bound is also satisfied for weakly interacting theories. [^36] For a strongly interacting example, see the case study of the $\mathcal{N}=4$ super Yang-Mills in the introduction. In general, our constraints (Eq. (4.38), or equivalently, Eq. (4.39)) can be explored in principle directly using CFT techniques. Finding a general proof would be highly non-trivial evidence for the rQFC. We leave this investigation to future work.
Note that since we are free to choose the operators $O^{(q)}_{f}$ which creates our near-vacuum state, if our CFT contains superselection sectors (say associated to some global symmetry), then by picking sector-specific $O^{(q)}_{f}$ we obtain our bound within each superselection sector. Relatedly, we have so far used a compact smearing function $f$ which inevitably mixes sectors of different transverse spin in the theory. It would, however, be interesting to generalize our argument to situations where $O^{(q)}_{f}$ can pick out specific transverse sector $j$ contributions of the $\mathcal{E}\times\mathcal{E}$ OPE. This would lead to the following stronger bounds on the spectrum:
$\delta_{j,1}\leq 2(d-1),~~~\text{for each even $j$},$ (4.41)
or equivalently, the coincident limit (4.39) is non-zero for each superselection sector.
Let us end by suggesting a more universal bound which we speculate might be true in all QFTs and implies the above-mentioned bounds. Roughly speaking, our bound (4.38) comes from the rQFC enforcing that the magnitude of (4.36) not go to zero as $\Sigma\to 0$ by a larger power than $\Sigma^{d-2}$, whenever $\delta\tilde{S}^{(1)}_{\rm ren}/\delta\tilde{V}(\tilde{y}^{a})\neq 0$. We have made two simplifying assumptions to extract this bound. We restricted to near vacuum-states, and, in particular, to ones in which $\langle T_{ij}\rangle=0$ in a neighborhood of the deformation by picking our smearing function $f$ to have no support there. Certainly, one can consider states with the latter property that are not near-vacuum, e.g., by setting $\epsilon=1$ in unitaries (4.4). We then still expect that the above-mentioned bound on the magnitude of (4.36) to hold. The only difference is that we cannot as directly connect it to the lightray OPE. Perhaps, the restriction to $\langle T_{ij}\rangle=0$ at the locus of null deformation is not essential for the argument either, and is only a convenient choice for the computation. We therefore speculate here that the following universal bound holds in the Rindler wedge for all QFTs:
@@
When applied to our near-vacuum states (4.4) in a CFT, the bound (4.42) reduces to our $\delta_{0}\leq 2(d-1)$ bound, though it is in general stronger. Note also that Eq. (4.42) is indeed stronger than the QNEC. This is because it is understood that the QNEC for a deformation family $\tilde{V}^{(1)}_{\lambda}(\tilde{y}^{a})$ implies the QNEC for a deformation family $\tilde{V}^{(2)}_{\lambda}(\tilde{y}^{a})$ if $\tilde{V}^{(1)}_{\lambda}(\tilde{y}^{a})<\tilde{V}^{(2)}_{\lambda}(\tilde{y}^{a})$, because the difference between the two has a definite sign due to strong subadditivity of von Neumann entropy. Therefore, if something stronger than the QNEC, such as Eq. (4.42) holds for deformation profiles in the limit of very narrow width, the QNEC for any deformation follows. We leave further exploration of Eq. (4.42) for future work.
@@ Acknowledgements
We are grateful to Ido Ben-Dayan, Raphael Bousso, Tom Faulkner, Luca Iliesiu, Adam Levine, Raghu Mahajan, Juan Maldacena, Hervé Partouche, Geoff Penington, David Simmons-Duffin, Zhenbin Yang, Aron Wall, and Alexander Zhiboedov for insightful discussions. We especially thank David Simmons-Duffin and Alexander Zhiboedov for very helpful exchanges about lightray operators, and Raghu Mahajan for comments on an earlier draft.

### v2

@@ 3 The rQFC and QFC in JT gravity coupled to a QFT
$\left(\frac{\Sigma}{L}\right)^{d-2-\delta}\gtrsim\left(\frac{\ell_{S}}{L}\right)^{d-2}.$ (4.35)
Therefore, the rQFC implies that $\delta\geq 0$. We can summarize this bound as:
$\left.\frac{1}{\sqrt{h_{\lambda}}}\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}=O\left(\frac{\epsilon}{L^{d-1}}\right)\implies\frac{1}{\sqrt{\tilde{h}_{\lambda}}}\partial_{\lambda}\left.\frac{\delta\tilde{S}^{(1)}_{\rm ren}}{\delta\tilde{V}(\tilde{y}^{a})}\right\rvert_{\lambda=0}\leq O(\frac{\epsilon^{2}\Sigma^{d-2}}{L^{2d-2}}).$ (4.36)
By the expression on the RHS, we mean that the second derivative of the von Neumann entropy cannot approach zero faster than $\Sigma^{d-2}$ as $\Sigma\to 0$. This bound is stronger than the QNEC in this class of states, since the QNEC merely requires that the RHS of (4.36) is non-positive.
Let us end by suggesting a more universal bound which we speculate might be true in all QFTs and implies the above-mentioned bounds. Roughly speaking, our bound (4.36) comes from the rQFC enforcing that the QNEC does not get close to saturation as $\Sigma\to 0$ by a larger power than $\Sigma^{d-2}$, whenever $\delta\tilde{S}^{(1)}_{\rm ren}/\delta\tilde{V}(\tilde{y}^{a})\neq 0$ at the relevant order in $\epsilon$. We have made two simplifying assumptions to extract this bound. We restricted to near vacuum-states, and, in particular, to ones in which $\langle T_{ij}\rangle=0$ in a neighborhood of the deformation by picking our smearing function $f$ to have no support there. Certainly, one can consider states with the latter property that are not near-vacuum, e.g., by setting $\epsilon=1$ in unitaries (4.4). We then still expect some verison of the above-mentioned bound on the magnitude of (4.33) to hold. The only difference is that we cannot as directly connect it to the product of $\mathcal{E}$ operators. Perhaps, the restriction to $\langle T_{ij}\rangle=0$ at the locus of null deformation is not essential for the argument either, and is only a convenient choice for the computation. We therefore speculate here that the following universal bound holds in the Rindler wedge for all QFTs:
@@
When applied to our near-vacuum states (4.4) in a CFT, the bound (4.37) implies Eq. (4.36). We leave further exploration of it to future work.
@@ Acknowledgements
We are grateful to Ido Ben-Dayan, Raphael Bousso, Tom Faulkner, Luca Iliesiu, Adam Levine, Raghu Mahajan, Juan Maldacena, Hervé Partouche, Geoff Penington, David Simmons-Duffin, Zhenbin Yang, Aron Wall, and Alexander Zhiboedov for insightful discussions. We especially thank David Simmons-Duffin and Alexander Zhiboedov for very helpful exchanges about lightray operators, Aron Wall for comments on an earlier draft, and Raghu Mahajan for comments on an earlier draft, and for pointing out an error in the first version of this paper.
@@ Appendix A JT gravity as dimensional reduction
Appendix D The $\mathcal{E}\times\mathcal{E}$ OPE
As discussed in Sec. 4, the rQFC implies that $\delta\geq 0$ in the following expression:
$\int_{-\infty}^{\infty}ds~e^{2}\langle\tilde{O}^{(1)}_{f}\mathcal{E}(\tilde{y}^{\prime a})\mathcal{E}(\tilde{y}^{a})\tilde{O}^{(1)}_{f_{s}}\rangle\stackrel{{\scriptstyle\tilde{y}^{\prime}\to\tilde{y}}}{{=}}O(|\tilde{y}-\tilde{y}^{\prime}|^{{-\delta}})$ (D.1)
The coincident limit $\tilde{y}^{\prime}\to\tilde{y}$ is in particular controlled by the OPE of two $\mathcal{E}$’s. This is a less conventional OPE which involves light-ray operators (i.e., operators with support on null generators of the $\tilde{u}=0$ plane). See [73, 74] for a proper discussion of light-ray operators and their OPEs.
[…]
$\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\stackrel{{\scriptstyle|\tilde{y}|\to 0}}{{=}}$ $\sum_{j,n}c_{j,n}~|\tilde{y}|^{\delta_{j,n}-2(d-1)}\hat{y}^{a_{1}}\cdots\hat{y}^{a_{j}}~\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}(0)+\text{descendents of $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$}$, (D.2)
where $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$ denotes a so-called light-ray operator with scaling dimension $\delta_{j,n}$, and transverse spin $j$ (see Fig. 12). Each operator $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$ is a non-local operator, and could be defined through smearing a pair local CFT operators in a certain way along a null geodesic. We will not cover the details of the light-ray OPE story here, but its structure (D.2) is sufficient for us to extract a certain bound on the spectrum $\delta_{j,n}$. Conveniently, the spectrum $\delta_{j,n}$ could be described simply in terms of the spectrum of local CFT data, allowing us the ability to phrase the bound in terms of data of the local CFT spectrum. Let $J$ denote total spin (as opposed to transverse spin $j$), and let $\Delta^{\text{even}}_{j,n}(J)$ be the analytically continued in $J$ of the scaling dimension of the n-th Regge trajectory in the transverse spin $j$ sector. We then have [75, 73, 74]:
$\delta_{j,n}=$ $\Delta^{\text{even}}_{j,n}(J=3)-1,~~~j=0,2,4.$ (D.3)
$\delta_{j,n}=$ $\Delta^{\text{even}}_{j=4,n}(J=3+2p)-1,~~~j=4+2p,~(p\in\mathbb{Z}^{+}).$ (D.4)
It is easy to see from the OPE (D.2), that the leading $\Sigma\to 0$ behavior of Eq. (4.27) is given by the smallest $\delta_{j,n}$ which is contributing in the state. A particularly drastic way for the rQFC implication that $\delta\geq 0$ in Eq. (D.1) to be violated is if the quantity $\delta_{0}$ defined as
$\delta_{0}\equiv\text{min}_{n,j}\delta_{j,n}.$ (D.5)
is such that in any state we obtain $\delta<0$. It is easy to see that for this not to happen we must have:
$\delta_{0}\leq 2(d-1).$ (D.6)
In fact, Eq. (D.6) is equivalent to stating that:
$\lim_{|\tilde{y}|\to 0}\mathcal{E}(0)\mathcal{E}(\tilde{y}^{a})\neq 0,$ (D.7)
Note that this is a much weaker constraint than the one discussed in the body of the paper. In fact, it trivially follows from demanding that $\mathcal{E}$ is non-zero as an operator. [^38] We thank Raghu Mahajan for pointing this out..
Figure 12: The OPE of two averaged null energy operators $\mathcal{E}(\tilde{y}^{a})=\int_{-\infty}^{\infty}d\lambda~T_{\tilde{v}\tilde{v}}(\tilde{v}=\lambda,\tilde{y}^{a})$ is shown. The RHS of the OPE consists of so-called light-ray operators $\mathbb{O}^{(j,n)}_{a_{1}\cdots a_{j}}$, labeled by a transverse spin $j$, and corresponding to a certain $n$-th (even spin) Regge trajectory. In our states, the second derivative of the von Neumann entropy under the null deformations of the Rindler wedge are controlled by this OPE: the more singular (as $|\tilde{y}|\to 0$) the OPE, the smaller the second derivative becomes in the limit of deformation width going to zero.
Let us now investigate how close can $\delta_{0}$ get to 2(d-1) in a few simple cases. First, in the case of free theories, the leading Regge trajectory ($n=1$) satisfies:
$\Delta^{\text{even}}_{j=0,1}(J)=d-2+J.$ (D.8)
Therefore, from Eq. (D.4) we have $\delta_{j=0,n=1}=d$, which satisfies the bound (D.6) with room to spare in $d\geq 3$.
Interestingly, we can obtain $\delta_{0}=2(d-1)$ in planar $\mathcal{N}=4$ SU(N)(N) super Yang-Mills in the large t’Hooft coupling limit. [^39] We thank David Simmons-Duffin for explaining to us the story of $\mathcal{N}=4$ Super Yang-Mills. In the $N\to\infty$, limit, we can delineate the Regge trajectories of single and double-trace operators. It turns out that in the large t’Hooft coupling limit, the single-trace $\text{min}_{j}\Delta^{\text{even}}_{j,1}(J)$ can be arbitrarily increased for each $J$, but the leading double-trace trajectory satisfies $\text{min}_{j}\Delta^{\text{even}}_{j,1}(J)=4+J$ (approaching it from below in the $N\to\infty$ limit) [76, 73]. Therefore, the double-trace trajectory leads to $\text{min}_{j}\delta_{j,1}=6$, which saturates the bound (D.6) given d=4d=4.404040For instance, one can consider the trajectory associated with the stress-tensor double-twist $TT]_{J}$, which at large spin have twist (i.e., scaling dimension minus spin) $\tau(J)=2d+J-(J+4)=2d-4$. We then obtain $\Delta(J=3)=\tau(J=3)+3=2d-1$.